OpenStudy (anonymous):

The roots of the quadratic equation 2x^2+11x-40=0 are 2p and (q-2).Calculate the possible values of p and q.

OpenStudy (anonymous):

@rational

OpenStudy (rational):

start by finding the roots of given quadratic

OpenStudy (rational):

see if it factors

OpenStudy (anonymous):

\[2p+q-2=-\frac{ 11 }{ 2 }\]\[2p(q-2)=-20\]

OpenStudy (rational):

Ahh nice you want to use sum of roots and product of roots

OpenStudy (rational):

maybe plugin the value of "q-2" from first equation into second equation

OpenStudy (anonymous):

\[2p(-\frac{ 11 }{ 2 })=-20\] like this? @rational

OpenStudy (rational):

from the first equation we have \[ q-2 = -\frac{11}{2}-2p\] plugging this in second equation gives \[2p(-\frac{11}{2}-2p) = -20\] right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[-11p-4p^2=-20\]

OpenStudy (anonymous):

\[-4p^2-11p+20=0\]\[4x^2+11x-20=0\]

OpenStudy (anonymous):

p=5/4,-4

OpenStudy (anonymous):

q=9/2,-6

OpenStudy (anonymous):

Thnx @rational :)

OpenStudy (rational):

looks perfect! yw :)

OpenStudy (jhannybean):

Ahh this reminds me of Vieta's Formula :)